OpenStudy (anonymous):

Each number from 1 to 60 is written on one of 60 index cards. the cards are arranged in rows with equal lengths and no cards are left over. The sum of of the numbers in each row is 305. How many rows are there?

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Total sum of 1~60 = 60*61/2=1830 # of Rows = r 1830=305*r r=1830/305=6 There are 6 rows.

OpenStudy (anonymous):

erm.. wasn't the formula (first term+second term)*n/2=Sn?

OpenStudy (anonymous):

No it is (First Term + Last Term)*n/2=Sn

OpenStudy (anonymous):

Oh yes, but you didn't multiply n

OpenStudy (anonymous):

No, in this case, first term is 1 and last term is 60. So it is (1+60)*60/2 which is 61*60/2

OpenStudy (anonymous):

ahh ty